OpenStudy (anonymous):

sqrt(2x+1)-sqrt(2x-1)=1

OpenStudy (zpupster):

you need to square both sides (sqrt(2x+1)-sqrt(2x-1))^2=1^2 or (sqrt(2x+1)-sqrt(2x-1) (sqrt(2x+1)-sqrt(2x-1)=1 use foil F=(sqrt(2x+1)(sqrt(2x+1)=2x+1 O=(sqrt(2x-1)(sqrt(2x+1) I=-(sqrt(2x-1)(sqrt(2x+1) L=-sqrt(2x-1) -sqrt(2x-1) =2x-1 combine like terms 4x-2(sqrt(2x-1)(sqrt(2x+1) 4x-2sqrt(4x^2-1)=1 isolate the sqrt root -2sqrt(4x^2-1)=1-4x square both sides again 4(4x^2)-1=(1-4x)^2

OpenStudy (zpupster):

now multiply out the square on right and multiply terms on left you should get it from here if not ask

OpenStudy (anonymous):

Got it thanks!